# Matrix3D: Large Photogrammetry Model All-in-One

### v1

@@ 2 Related Work
Multi-view Stereo (MVS) builds upon the camera poses obtained by SfM to create dense 3D geometry. Traditional MVS methods [37, 88, 27, 26] depend on hand-crafted features and engineered regularizations to build dense correspondences and recover 3D points [48, 25], volumes [46, 90, 25], or depth maps [8, 102, 27]. Learning-based methods, including MVSNet [127] and follow-up works [128, 14, 31, 134], offer more powerful reconstruction with improved completeness and generalization capabilities. Typically, these methods assume well-calibrated camera parameters, which limits their robustness and applicability in real-world scenarios where calibration may be imprecise or unavailable.
@@
Masked Learning has achieved significant success in pre-training tasks for NLP [44, 7, 74, 75] and computer vision [36, 2, 50]. These methods have been proven to capture high-level semantics by masking parts of input and training models to reconstruct the masked contents. Recent research has extended this idea to multi-view image settings [113, 114], demonstrating improvements in downstream tasks such as optical flow and stereo matching. In this work, we further apply masked learning to multi-view and multi-modal training to develop an all-in-one photogrammetry model.
@@ 3 Method
In this section, we introduce Matrix3D, an all-in-one photogrammetry model for unified 3D reconstruction and generation. Matrix3D is built upon a multi-modal diffusion transformer. In the following, we describe the details of the framework design (Section 3.1), masking strategies (Section 3.2), dataset preparation (Section 3.3), training setup (Section 3.4), and downstream tasks (Section 3.5).
@@ 3.3 Dataset Preparation
Incomplete depth. Real-world datasets often provide incomplete depth. A default solution to use these data is discarding patches with invalid pixels, which results in low data utilization. Instead, we concatenate valid masks to all the depth maps for both condition maps and noisy maps, and let the model learn how to discard invalid pixels. This also allow us to utilize incomplete or even sparse depth input during inference.
@@ 3.4 Training Setup
We initialize our model with pre-trained Hunyuan-DiT checkpoints [51]. In the first stage, we train the model for 180K steps with a learning rate of 1e-4 on 64 NVIDIA A100 80G GPUs. In the latter two stages, we train the model for 20K steps for each stage with a learning rate of 1e-5 on 128 GPUs. The entire training takes approximately 20 days to complete.
@@ 4.1 Pose Estimation
Table 2: Novel view synthesis evaluation on GSO. Results are grouped by different view settings because some methods uses fixed novel poses. The best results are in bold and the second bests are underlined. $\phi$ and $\theta$ denote azimuth and elevation angles, respectively.
@@ 4.2 Novel View Synthesis
In this section, we benchmark novel view synthesis task on diffusion samples on GSO [23] dataset against prior multi-view diffusion methods, including Zero123 [59], Zero123XL [20], SyncDreamer [60], Wonder3D [61], and InstantMesh [119]. For methods generating novel views at fixed camera poses, we follow their view configuration and use different rendered ground truths when comparing with them. For methods allowing arbitrary view synthesis, we render 32 random viewpoints for evaluation. PSNR, SSIM, and LPIPS are adopted as evaluation metrics. Figure 4 demonstrates qualitative results on the CO3D and the ARKitScenes dataset. The quantitative results are shown in Table 2. Our method achieves the best results for most of the metrics.
@@ 4.3 Depth Prediction
Figure 6: Unposed sparse-view 3D reconstruction results.
4.4 3D Reconstruction
Monocular. In the following, we evaluate the 3D reconstruction performance from single images. Specifically, we compare Matrix3D with diffusion-based optimization methods including ImageDream [109], One2345++ [57], IM-3D [67], and CAT3D [29] in terms of CLIP scores following CAT3D [29]. Figure 5 and Table 6 illustrate the comparisons. Our method achieves comparable results to SOTA methods.
@@ 5 Conclusion
In this paper, we introduced Matrix3D, a unified model that effectively addresses multiple photogrammetry tasks including pose estimation, depth prediction, and novel view synthesis using a multi-modal diffusion transformer (DiT). By employing a mask learning strategy, Matrix3D supports flexible input/output combinations, and maximizes training data from incomplete datasets. Through multi-round, multi-view, multi-model interactive generation, users can perform single or few-shot generation with one single model. Extensive experiments show that Matrix3D achieves state-of-the-art performance in each downstream tasks mentioned above, demonstrating its versatility on photogrammetry applications.
@@ 7 Dataset Pre-processing
For the Objaverse dataset, we render all models into RGB and depth images for training. Specifically, each 3D object is first normalized at the world center within a bounding box of $[-1,1]^{3}$, and we render the whole scene from 32 random viewpoints. The render camera FoV is set to 50°$\degree$. The azimuth and elevation angles are randomly sampled in $[0\degree,360\degree]$ and $[-45\degree,90\degree]$. The camera distance to the world center is randomly sampled in [1.1, 1.6], and the height on the z-axis is set in [-0.4, 1.2]. We use a composition of random lighting from area lighting, sun lighting, point lighting, and spot lighting.
@@ 9.5 Limitation
Methods GT GT GT Align DTU ETH3D T&T
DeepV2D $\times$ $\times$ $\checkmark$ med 7.7 33.0 11.8 29.3 8.9 46.4
DUSt3R $\times$ $\times$ $\times$ med 2.76 77.32 4.71 61.74 5.54 56.38
Ours $\times$ $\times$ $\checkmark$ med 1.85 85.46 7.83 38.80 6.16 49.43
Table 10: Unposed MVD evaluation on DTU, ETH3D, and T&T.

### v2

@@ 2 Related Work
Multi-view Stereo (MVS) builds upon the camera poses obtained by SfM to create dense 3D geometry. Traditional MVS methods [37, 88, 27, 26] depend on hand-crafted features and engineered regularizations to build dense correspondences and recover 3D points [48, 25], volumes [46, 90, 25], or depth maps [8, 102, 27]. Learning-based methods [127, 128, 14, 31, 134] offer more powerful reconstruction with improved completeness and generalization capabilities. Typically, these methods assume well-calibrated camera parameters, which limits their robustness and applicability in real-world scenarios where calibration may be imprecise or unavailable.
@@
Masked Learning has achieved significant success in pre-training tasks for NLP [44, 7, 74, 75] and computer vision [36, 2, 50]. These methods are proven to capture high-level semantics by masking parts of input and training models to reconstruct the masked contents. Recent research has extended this idea to multi-view image settings [113, 114], demonstrating improvements in downstream tasks like optical flow and stereo matching. In this work, we further apply masked learning to multi-view and multi-modal training to develop an all-in-one photogrammetry model.
@@ 3 Method
In this section, we introduce Matrix3D, an all-in-one photogrammetry model for unified 3D reconstruction and generation. In the following, we describe the details of the framework design (Section 3.1), masking strategies (Section 3.2), dataset preparation (Section 3.3), training setup (Section 3.4), and downstream tasks (Section 3.5).
@@ 3.3 Dataset Preparation
Normalization. Due to the highly diverse distributions of existing datasets, including variations in scale and scene type, preprocessing them consistently poses a challenge. To address this, we apply scene normalization and camera normalization. Please check the supplementary for details.
Incomplete depth. Real-world datasets often provide incomplete depth. A default solution to use these data is discarding patches with invalid pixels, which results in low data utilization. Instead, we concatenate valid masks to all the depth maps for both condition maps and noisy maps, and let the model learn how to discard invalid pixels. This also allow us to utilize sparse depth input during inference.
@@ 3.4 Training Setup
We initialize our model with pre-trained Hunyuan-DiT checkpoints [51]. In the first stage, we train the model for 180K steps with a learning rate of 1e-4 on 64 NVIDIA A100 80G GPUs. In the latter two stages, we train the model for 20K steps for each stage with a learning rate of 1e-5 on 128 GPUs. The entire training takes approximately 20 days.
@@ 4.1 Pose Estimation
Table 2: Novel view synthesis (diffusion samples) evaluation on GSO. Results are grouped by different view settings because some methods uses fixed poses. Best results are in bold and the second best are underlined. $\phi$ and $\theta$ denote azimuth and elevation angles.
@@ 4.2 Novel View Synthesis
In this section, we benchmark novel view synthesis task on diffusion samples on GSO [23] dataset against prior multi-view diffusion methods, including Zero123 [59], Zero123XL [20], SyncDreamer [60], Wonder3D [61], and InstantMesh [119]. For methods generating novel views at fixed camera poses, we follow their view configuration and use different rendered ground truths when comparing with them. For methods allowing arbitrary view synthesis, we render 32 random viewpoints for evaluation. PSNR, SSIM, and LPIPS are adopted as evaluation metrics. Figure 4 shows qualitative results on CO3D and ARKitScenes dataset. The quantitative results are shown in Table 2. Our method achieves the best results for most of the metrics.
4.3 Depth Prediction
@@
Figure 7: Unposed sparse-view 3D reconstruction results.
4.4 3D Reconstruction
Monocular. In the following, we evaluate the 3D reconstruction performance from single images. Specifically, we compare Matrix3D with diffusion-based optimization methods including ImageDream [109], One2345++ [57], IM-3D [67], and CAT3D [29] in terms of CLIP scores following CAT3D. Figure 5 and Table 6 illustrate the comparisons. Our method achieves comparable results to SOTA methods.
@@ 5 Conclusion
In this paper, we introduced Matrix3D, a unified model that effectively addresses multiple photogrammetry tasks including pose estimation, depth prediction, and novel view synthesis using a multi-modal diffusion transformer (DiT). By employing a mask learning strategy, Matrix3D supports flexible input/output combinations, and maximizes training data from incomplete datasets. Through multi-round, multi-view, multi-model interactive generation, users can perform single or few-shot generation with one single model. Extensive experiments show that Matrix3D achieves SOTA performance in pose estimation and novel view synthesis tasks, showing its versatility on photogrammetry applications.
6 Acknowledgments
This work is supported by National Natural Science Foundation of China (62472213, 62025108), Gusu Innovation & Entrepreneurship Leading Talents Program (ZXL2024361), and Hong Kong RGC GRF 16206722.
@@ 8 Dataset Pre-processing
Objaverse Rendering. For the Objaverse dataset, we render all models into RGB and depth images for training. Specifically, each 3D object is first normalized at the world center within a bounding box of $[-1,1]^{3}$, and we render the whole scene from 32 random viewpoints. The render camera FoV is set to 50°$\degree$. The azimuth and elevation angles are randomly sampled in $[0\degree,360\degree]$ and $[-45\degree,90\degree]$. The camera distance to the world center is randomly sampled in [1.1, 1.6], and the height on the z-axis is set in [-0.4, 1.2]. We use a composition of random lighting from area lighting, sun lighting, point lighting, and spot lighting.
@@ 10.2 Point Cloud Fusion
Methods GT GT GT Align DTU ETH3D T&T
DeepV2D $\times$ $\times$ $\checkmark$ med 7.7 33.0 11.8 29.3 8.9 46.4
DUSt3R $\times$ $\times$ $\times$ med 2.76 77.32 4.71 61.74 5.54 56.38
Ours $\times$ $\times$ $\checkmark$ med 1.85 85.46 7.83 38.80 6.16 49.43
Table 9: Unposed MVD evaluation on DTU, ETH3D, and T&T.